# Descending finite projective modules from a Novikov ring

### v1

@@ 1. Introduction
$A(\!(t^{\Gamma})\!)\rightrightarrows A(\!(t^{\Gamma},u^{\Gamma})\!)\rightrightarrows\to A(\!(t^{\Gamma},u^{\Gamma},v^{\Gamma})\!)$
@@
In general, the ring $A(\!(t^{\Gamma},u^{\Gamma})\!)$ is much larger than the tensor product $A(\!(t^{\Gamma})\!)\otimes_{A}A(\!(t^{\Gamma})\!)$, which is usually where the isomorphism of the descent datum is defined. As a consequence, standard descent results for finite projective modules are inapplicable to this setting.
The theorem is motivated by question in perfectoid geometry, namely that of descending a vector bundle from $\operatorname{Spd}A$ to a vector bundle on $\operatorname{Spec}A$, where $A$ is a discrete perfect ring in characteristic $p$. Using Theorem 4.9 with $\Gamma=\mathbb{Z}[p^{-1}]$, we deduce the following result as an immediate consequence.
Theorem 1.3 (Corollary 5.2, 5.4).
@@
{vector bundles on $\{\text{vector bundles on }\operatorname{Spec}A\}\to\{\text{vector bundles on }\operatorname{Spd}A\}$
@@
It is not true in general that a v-sheaf étale over $\operatorname{Spd}A$ descends to a scheme étale over $\operatorname{Spec}A$. For example, consider the perfect $\mathbb{F}_{p}$-algebra $A=\mathbb{F}_{p}[t^{1/p^{\infty}}]$. The associated v-sheaf is $(\operatorname{Spd}A)(R,R^{+})=R^{+}$, and admits a v-subsheaf $X(R,R^{+})=R^{\circ\circ}$. Then $X\to\operatorname{Spd}A$ is an open embedding, hence étale, but does not correspond to an étale $A$-scheme. Indeed, for every closed point $x\in\operatorname{Spec}A$, the fiber $\operatorname{Spd}k(x)\times_{\operatorname{Spd}A}X$ is empty when $x\neq 0$ and is $\operatorname{Spd}\mathbb{F}_{p}$ when $x=0$.
@@
The second part of the theorem, Corollary 5.4, will be used in a future work [DKvHZ24] to construct integral models of Hodge type Shimura varieties at parahoric level. Specifically, it allows one to construct a finite étale cover of the special fiber a Shimura variety from the finite étale cover of the associated v-sheaf.
@@
Ian Gleason announced via private communication an alternative proof of Corollary 5.4, which works by comparing the category of étale $\mathbb{F}_{\ell}$-local systems on $\operatorname{Spd}A$ and $\operatorname{Spec}A$.
@@ 1.1. Outline of proof
We sketch the main ideas that go into the proof of Theorem 4.9. In Section 3, we introduce the notion of a Novikov isocrystal, which somewhat easier to control than a descent datum as it involves only one variable. In Subsection 4.1, we construct a functor from the category of Novikov descent data for $\Gamma=\mathbb{R}$ to the category of Novikov isocrystals, and prove that this functor is always fully faithful. This allows us to reduce the effectivity of a Novikov descent datum into a similar statement about the associated Novikov isocrystal.
In Subsection 4.2, we prove effectivity of Novikov descent data when $A$ is reduced and $\Gamma=\mathbb{R}$. We first take care the case when $A$ is an algebraically closed field by using the structure theory of Novikov isocrystals developed in Section 3. Next, we check that when $A$ is an arbitrary product of algebraically closed fields, effectivity over $A$ follows from that over each field. This allows to deduce effectivity for a general reduced ring $A$ by embedding $A$ into the product of the algebraic closures of its residue fields.
@@ 1.2. Acknowledgements
I would like to thank Patrick Daniels, Pol van Hoften, and Mingjia Zhang for help formulating the problem, which arose in our work on integral models of Shimura varieties, as well as insightful conversations. I would also like to thank my advisor Richard Taylor for his encouragements and diligence during our meetings, where we carefully went through parts of the argument. I also would like to thank Brian Conrad, Ian Gleason, and Kiran Kedlaya for helpful comments and suggestions.
@@ 2. Novikov rings
We note that the functor $M\mapsto M(\!(t^{\Gamma})\!)$ sending $A$-modules to $A(\!(t^{\Gamma})\!)$-modules commutes with limits and finite colimits. In particular, when $M$ is a finitely presented $A$-module, the natural map
$M\otimes_{A}A(\!(t^{\Gamma})\!)\to M(\!(t^{\Gamma})\!)$
is an isomorphism. As a consequence, for any other $A$-module $N$ there is a canonical isomorphism
$\operatorname{Hom}_{A(\!(t^{\Gamma})\!)}(M(\!(t^{\Gamma})\!),N(\!(t^{\Gamma})\!))\cong\operatorname{Hom}_{A}(M,N)(\!(t^{\Gamma})\!)$
of $A(\!(t^{\Gamma})\!)$-modules.
@@
$\pi_{1}^{\ast}$ $\colon A(\!(t^{\Gamma})\!)\to A(\!(t^{\Gamma},u^{\Gamma})\!);$ $t$ t,maps-$\mapsto t,$
@@
$A\to A(\!(t^{\Gamma})\!)\rightrightarrows A(\!(t^{\Gamma},u^{\Gamma})\!)\rightrightarrows\to A(\!(t^{\Gamma},u^{\Gamma},v^{\Gamma})\!).$
We will also use the notation $\rho_{1}^{\ast}=\pi_{12}^{\ast}\circ\pi_{1}^{\ast}=\pi_{13}^{\ast}\circ\pi_{1}^{\ast}$ and similarly $\rho_{2}^{\ast}=\pi_{12}^{\ast}\circ\pi_{2}^{\ast}=\pi_{23}^{\ast}\circ\pi_{1}^{\ast}$ and $\rho_{3}^{\ast}=\pi_{13}^{\ast}\circ\pi_{2}^{\ast}=\pi_{23}^{\ast}\circ\pi_{2}^{\ast}$. For a ring homomorphism $f^{\ast}\colon A\to B$ and an $A$-module $M$, we will also denote $f^{\ast}M=M\otimes_{A,f^{\ast}}B$, and abuse notation to write $f^{\ast}\colon M\to f^{\ast}M$ for the map $m\mapsto m\otimes 1$.
@@
is a $A(\!(t^{\Gamma},u^{\Gamma})\!)$-linear automorphism satisfying the cocycle condition
[…]
We denote by $\mathsf{NovDD}(A,\Gamma)$ the category of Novikov descent data.
@@
$\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\Gamma);\quad M_{0}\mapsto(M_{0}\otimes_{A}A(\!(t^{\Gamma})\!),\operatorname{id}_{M_{0}\otimes_{A}A(\!(t^{\Gamma},u^{\Gamma})\!)}),$
which will turn out to be an equivalence by Theorem 4.9.
@@
where $a_{i}\neq 0$ for all $i\in I$ and $(d_{i,1},\dotsc,d_{i,n})\in\Gamma^{n}$ are all distinct, we define the support of $x$ as
@@
The support of $0\in A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ is empty. We always have $\operatorname{supp}(x+y)\subseteq\operatorname{supp}(x)\cup\operatorname{supp}(y)$ and $\operatorname{supp}(xy)\subseteq\operatorname{supp}(x)+\operatorname{supp}(y)$, where $+$ is the Minkowski sum defined as $X+Y=\{x+y:x\in X,y\in Y\}$. By definition, for every element $x\in A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ or $x\in M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$, the intersection of $\operatorname{supp}(x)$ with the half-space
@@
be a $A(\!(t^{\Gamma})\!)$-linear map compatible with the descent data on both sides. Viewing $f$ as an element of $\operatorname{Hom}_{A}(M_{0},N_{0})(\!(t^{\Gamma})\!)$, we see that compatibility with descent data translates to
@@ 2.1. Topological rings and modules
$t^{N}M(\!(t^{\mathbb{R}})\!)\subseteq\sum_{i=1}^{n}x_{i}A[[t^{\mathbb{R}}]].$
In particular, for every $m\in M$, the element $t^{N}m\in M(\!(t^{\mathbb{R}})\!)$ can be written as $A[[t^{\mathbb{R}}]]$-linear combination of $x_{1},\dotsc,x_{n}$. Therefore once we write $x_{i}=\sum_{j}m_{ij}t^{d_{ij}}$, the collection of $m_{ij}$ with $d_{ij}\leq N$ generate $M$, and so $M$ is finitely generated.

### v2

@@ 1. Introduction
$A(\!(t^{\Gamma})\!)\rightrightarrows A(\!(t^{\Gamma},u^{\Gamma})\!)\raisebox{1.70717pt}{$\rightrightarrows$}\hskip-10.00002pt\raisebox{-1.70717pt}{$\rightrightarrows$}A(\!(t^{\Gamma},u^{\Gamma},v^{\Gamma})\!)$
@@
In general, the ring $A(\!(t^{\Gamma},u^{\Gamma})\!)$ is much larger than the tensor product $A(\!(t^{\Gamma})\!)\otimes_{A}A(\!(t^{\Gamma})\!)$, which is usually where the isomorphism of the descent datum is defined over. As a consequence, standard descent results for finite projective modules are inapplicable to this setting.
Remark 1.3.
This may heuristically be interpreted as an analytic-to-discrete descent result for adic spaces. For example, let $k$ be a field, consider the field k((t))k(\!(t)\!) of Laurent series with the $t$-adic topology, and take its adic spectrum $X=\operatorname{Spa}(k(\!(t)\!),k[[t]])$. Then we have
$Y=X\times_{\operatorname{Spa}(k,k)}X=\bigcup_{n\geq 1}\operatorname{Spa}\Bigl(k\Bigl[t,u,\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr]^{\wedge}\Bigl[\frac{1}{tu}\Bigr],k\Bigl[t,u,\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr]^{\wedge}\Bigr)$
in the category of adic spaces, where we use the $(tu)$-adic topology for the completion. The adic space $Y$ can also be regarded as the punctured open unit disc over k((t))k(\!(t)\!), and we have $H^{0}(Y,\mathscr{O}_{Y})=k(\!(t,u)\!)$. Then Theorem 1.1 implies that the category of vector bundles on $X$ equipped with descent data over $Y$ is equivalent to the category of vector bundles on $\operatorname{Spa}(k,k)$.
The theorem is motivated by question in perfectoid geometry, namely that of descending a vector bundle from $\operatorname{Spd}A$ to a vector bundle on $\operatorname{Spec}A$, where $A$ is a discrete perfect ring in characteristic $p$. Using Theorem 1.1 with $\Gamma=\mathbb{Z}[p^{-1}]$, we deduce the following result as an immediate consequence. We refer the reader to Section 5 for the definitions of the objects involved.
Theorem 1.4 (Corollary 5.3, 5.5).
@@
$\{\text{vector bundles on }\operatorname{Spec}A\}\to\{\text{vector bundles on }\operatorname{Spd}A\}$
@@
It is not true in general that a v-sheaf étale over $\operatorname{Spd}A$ descends to a scheme étale over $\operatorname{Spec}A$. Consider the following example from [SW20, Example 18.2.1]. Take the perfect $\mathbb{F}_{p}$-algebra $A=\mathbb{F}_{p}[t^{1/p^{\infty}}]$, which has associated v-sheaf $(\operatorname{Spd}A)(R,R^{+})=R^{+}$ that admits a v-subsheaf $X(R,R^{+})=R^{\circ\circ}$. Then $X\to\operatorname{Spd}A$ is an open embedding, hence étale, but does not correspond to an étale $A$-scheme. Indeed, for every closed point $x\in\operatorname{Spec}A$, the fiber $\operatorname{Spd}k(x)\times_{\operatorname{Spd}A}X$ is empty when $x\neq 0$ and is $\operatorname{Spd}\mathbb{F}_{p}$ when $x=0$.
@@
The second part of Theorem 1.4 is used in [DKvHZ24] to construct integral models of Hodge type Shimura varieties at parahoric level. Specifically, it allows one to construct a finite étale cover of the special fiber of a Shimura variety from the finite étale cover of the associated v-sheaf.
@@
Ian Gleason announced via private communication an alternative proof of Corollary 5.5, i.e., the finite étale part of Theorem 1.4, which works by comparing the category of étale $\mathbb{F}_{\ell}$-local systems on $\operatorname{Spd}A$ and $\operatorname{Spec}A$.
@@ 1.1. Outline of proof
We sketch the main ideas that go into the proof of Theorem 1.1. In Section 3, we introduce the notion of a Novikov isocrystal, which is somewhat easier to control than a descent datum as it involves only one variable. In Subsection 4.1, we construct a functor from the category of Novikov descent data for $\Gamma=\mathbb{R}$ to the category of Novikov isocrystals, and prove that this functor is always fully faithful. This allows us to reduce the effectivity of a Novikov descent datum into a similar statement about the associated Novikov isocrystal.
In Subsection 4.2, we prove effectivity of Novikov descent data when $A$ is reduced and $\Gamma=\mathbb{R}$. We first take care the case when $A$ is an algebraically closed field by using the structure theory of Novikov isocrystals developed in Section 3. Next, we check that when $A$ is an arbitrary product of algebraically closed fields, effectivity over $A$ from that over each field. This allows to deduce effectivity for a general reduced ring $A$ by embedding $A$ into the product of the algebraic closures of its residue fields.
@@ 1.2. Acknowledgements
I would like to thank Patrick Daniels, Pol van Hoften, and Mingjia Zhang for help formulating the problem, which arose in our work on integral models of Shimura varieties, as well as insightful conversations. I would also like to thank my advisor Richard Taylor for his encouragements and diligence during our meetings, where we carefully went through parts of the argument. I also would like to thank Brian Conrad, Ian Gleason, Kiran Kedlaya, and the anonymous referee for helpful comments and suggestions.
@@ 2. Novikov rings
We may also describe the ring $A(\!(t^{\Gamma})\!)$ and the module $M(\!(t^{\Gamma})\!)$ as follows. We first consider the polynomial ring $A[t^{\Gamma}]$ and the module $M[t^{\Gamma}]$ where elements of $A[t^{\Gamma}]$ are finite sums of monomials $a_{i}t^{d_{i}}$. We endow $A[t^{\Gamma}]$ with the topology where $\bigoplus_{c<d\in\Gamma}At^{d}\subseteq A[t^{\Gamma}]$ for $c\in\mathbb{R}$ form a neighborhood basis of 0, and similarly for $M[t^{\Gamma}]$. Then $A(\!(t^{\Gamma})\!)$ and $M(\!(t^{\Gamma})\!)$ are the completions of the topological ring $A[t^{\Gamma}]$ and the topological module $M[t^{\Gamma}]$.
Remark 2.3.
@@
Example 2.5.
When $\Gamma=\mathbb{Z}_{\leq 0}$, we have $A(\!(t^{\Gamma},u^{\Gamma})\!)=A[t^{-1},u^{-1}]$, and when $\Gamma=\mathbb{Z}_{\geq 0}$, we have $A(\!(t^{\Gamma},u^{\Gamma})\!)=A[[t,u]]$. When $\Gamma=\mathbb{Z}$, the ring A((t,u))A(\!(t,u)\!) is in general much larger than the ring $A[[t,u]]{}[(tu)^{-1}]$ since it contains convergent power series such as
$\sum_{n=0}^{\infty}t^{-n}u^{n^{2}}\in A(\!(t,u)\!).$
On the other hand, even if $A$ is a field, the ring A((t,u))A(\!(t,u)\!) is not a field because tut-u is not invertible.
Remark 2.6.
As in Remark 2.2, we may construct the Novikov rings $A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ as a completion of the polynomial ring $A[t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma}]$. Here, we given $A[t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma}]$ the topology generated by additive cosets of
$\bigoplus_{d_{1},\dotsc,d_{n}\in\Gamma,\sum_{j}s_{j}d_{j}\geq C}At_{1}^{d_{1}}\dotsm t_{n}^{d_{n}}$
where $s_{1},\dotsc,s_{n}\in\mathbb{R}_{>0}$ and $C\in\mathbb{R}$. This is indeed a topological ring, and its completion is $A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$. One can similarly define a topological $A[t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma}]$-module $M[t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma}]$ whose completion is $M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$. In particular, it follows from this discussion that $A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ indeed has a natural ring structure, and $M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ indeed has a natural structure of a module over it.
Remark 2.7.
In the case when $\Gamma\subseteq\mathbb{R}$ is a nontrivial subgroup, there is a more geometric interpretation. Fix an element $0<\gamma\in\Gamma$, and consider the ring $A[[t_{1}^{\Gamma_{\geq 0}},\dotsc,t_{n}^{\Gamma_{\geq 0}}]]$, which is completion of $A[t_{1}^{\Gamma_{\geq 0}},\dotsc,t_{n}^{\Gamma_{\geq 0}}]$ with respect to the $(t_{1}^{\gamma},\dotsc,t_{n}^{\gamma})$-adic topology. This defines an affinoid pre-adic space
$X=\operatorname{Spa}(A[[t_{1}^{\Gamma_{\geq 0}},\dotsc,t_{n}^{\Gamma_{\geq 0}}]],A[[t_{1}^{\Gamma_{\geq 0}},\dotsc,t_{n}^{\Gamma_{\geq 0}}]]).$
The locus on which $\lvert t_{1}\rvert,\dotsc,\lvert t_{n}\rvert>0$ is the union of the rational localizations
$\operatorname{Spa}(R_{m},R_{m}^{+})$ $=\{\lvert-\rvert\in X:\lvert t_{j}\rvert>\lvert t_{i}\rvert^{m}>0\text{ for all }1\leq i,j\leq n\},$
[…]
Lemma 2.8.
The functor from $A$-modules to $A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$-modules sending $M\mapsto M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ is exact.
Proof.
Given surjective $A$-module homomorphism $f\colon M_{1}\twoheadrightarrow M_{2}$, the induced map
$f(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)\colon M_{1}(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)\to M_{2}(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$
surjective because we can lift each nonzero coefficient individually. The kernel $\ker f(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ consists of those series whose coefficients all lie in $\ker f$, and hence is $(\ker f)(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$. ∎
In particular, when $M$ is a finitely presented $A$-module, the natural map
$M\otimes_{A}A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)\to M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$
is an isomorphism. As a consequence, for any other $A$-module NN there is a canonical isomorphism
$\operatorname{Hom}_{A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)}(M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!),N(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!))\cong\operatorname{Hom}_{A}(M,N)(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$
of $A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$-modules.
2.1. Novikov descent data
@@
$\pi_{1}^{\ast}$ $\colon A(\!(t^{\Gamma})\!)\to A(\!(t^{\Gamma},u^{\Gamma})\!);$ $t$ $\mapsto t,$
@@
$A\to A(\!(t^{\Gamma})\!)\rightrightarrows A(\!(t^{\Gamma},u^{\Gamma})\!)\raisebox{1.67007pt}{$\rightrightarrows$}\hskip-10.00002pt\raisebox{-1.67007pt}{$\rightrightarrows$}A(\!(t^{\Gamma},u^{\Gamma},v^{\Gamma})\!).$
We will also use the notation $\rho_{1}^{\ast}=\pi_{12}^{\ast}\circ\pi_{1}^{\ast}=\pi_{13}^{\ast}\circ\pi_{1}^{\ast}$ and similarly $\rho_{2}^{\ast}=\pi_{12}^{\ast}\circ\pi_{2}^{\ast}=\pi_{23}^{\ast}\circ\pi_{1}^{\ast}$ and $\rho_{3}^{\ast}=\pi_{13}^{\ast}\circ\pi_{2}^{\ast}=\pi_{23}^{\ast}\circ\pi_{2}^{\ast}$. For a ring homomorphism $f^{\ast}\colon A\to B$ and an $A$-module $M$, we will also denote $f^{\ast}M=M\otimes_{A,f^{\ast}}B$, and write $f^{\ast}\colon M\to f^{\ast}M$ for the map $m\mapsto m\otimes 1$ by making an abuse of notation.
@@
is an $A(\!(t^{\Gamma},u^{\Gamma})\!)$-linear automorphism satisfying the cocycle condition, which states that the diagram
$\rho_{1}^{\ast}M$$\rho_{3}^{\ast}M$$\rho_{2}^{\ast}M.$$\scriptstyle{\pi_{13}^{\ast}\varphi}$$\scriptstyle{\pi_{12}^{\ast}\varphi}$$\scriptstyle{\pi_{23}^{\ast}\varphi}$
commutes. We denote by $\mathsf{NovDD}(A,\Gamma)$ the category of Novikov descent data.
@@
$\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\Gamma);\quad M_{0}\mapsto(M_{0}\otimes_{A}A(\!(t^{\Gamma})\!),\operatorname{id}_{M_{0}\otimes_{A}A(\!(t^{\Gamma},u^{\Gamma})\!)}).$
Remark 2.10.
Our main theorem, Theorem 1.1, may now be reformulated to state that the functor $\mathsf{Vect}(A)\to\mathsf{NovDD}(A,\Gamma)$ is an equivalence of categories.
@@
where $a_{i}\neq 0$ for all $i\in I$ and $(d_{i,1},\dotsc,d_{i,n})\neq(d_{j,1},\dotsc,d_{j,n})$ for $i\neq j$ in $I$, we define the support of $x$ as
@@
The support of $0\in A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ is empty. We always have $\operatorname{supp}(x+y)\subseteq\operatorname{supp}(x)\cup\operatorname{supp}(y)$ and $\operatorname{supp}(xy)\subseteq\operatorname{supp}(x)+\operatorname{supp}(y)$, where ++ is the Minkowski sum defined as $X+Y=\{x+y:x\in X,y\in Y\}$. By definition, for every element $x\in A(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$ or $x\in M(\!(t_{1}^{\Gamma},\dotsc,t_{n}^{\Gamma})\!)$, the intersection of $\operatorname{supp}(x)$ with the half-space
@@
be an $A(\!(t^{\Gamma})\!)$-linear map compatible with the descent data on both sides. This means that the diagram
$M_{0}(\!(t^{\Gamma},u^{\Gamma})\!)$$N_{0}(\!(t^{\Gamma},u^{\Gamma})\!)$$M_{0}(\!(t^{\Gamma},u^{\Gamma})\!)$$N_{0}(\!(t^{\Gamma},u^{\Gamma})\!)$$\scriptstyle{\pi_{1}^{\ast}f}$$\scriptstyle{\pi_{2}^{\ast}f}$
commutes, where both $\pi_{1}^{\ast}f$ and $\pi_{2}^{\ast}f$ are $A(\!(t^{\Gamma},u^{\Gamma})\!)$-linear. Viewing $f$ as an element of $\operatorname{Hom}_{A}(M_{0},N_{0})(\!(t^{\Gamma})\!)$, we see that compatibility with descent data translates to
@@ 2.2. Topological rings and modules
$t^{N}M[[t^{\mathbb{R}}]]\subseteq\sum_{i=1}^{n}x_{i}A[[t^{\mathbb{R}}]].$
In particular, for every $m\in M$, the element $t^{N}m\in M[[t^{\mathbb{R}}]]$ can be written as $A[[t^{\mathbb{R}}]]$-linear combination of $x_{1},\dotsc,x_{n}$. Therefore once we write $x_{i}=\sum_{j}m_{ij}t^{d_{ij}}$, the collection of $m_{ij}$ with $d_{ij}\leq N$ generate $M$, and so $M$ is finitely generated.